# On the Robustness of Verbal Confidence of LLMs in Adversarial Attacks

### v1

@@ 3.3.2 ConfidenceTriggers-AutoDAN
As an alternative method for more naturalistic triggers, we propose a variant of ConfidenceTriggers based on AutoDAN [37] (ConfidenceTriggers-AutoDAN). AutoDAN is an algorithm for automatically generating stealthy jailbreak prompts through a hierarchical genetic algorithm. Featuring the same structure as ConfidenceTriggers, hierarchical perturbations are produced each generation at the sentence-level through synonym replacement based on scores from a momentum word score dictionary, and at the paragraph-level using GPT4 [45] to rephrase pairs of parent prompts for Cross-Over and Mutation. Full details and the algorithm are provided in Appendix D.
@@ 3.4 Design of Confidence Elicitation
Our confidence elicitation methods are based on the framework proposed in [66], where CEMs primarily consist of a prompting strategy (i.e., task instruction) for generating verbalized numeric confidence scores. CEMs can also include sampling strategies for generating multiple LLM responses, which are then aggregated to ideally produce a more representative confidence. We use the self-random sampling strategy, which samples $K$ different output sequences $\{\bm{\mathcal{Y}}_{1},\bm{\mathcal{Y}}_{2},\dots,\bm{\mathcal{Y}}_{K}\}$ using a specific temperature parameter, from which confidence estimates $\{\mathcal{C}_{1},\dots\mathcal{C}_{K}\}$ are derived. To aggregate these disparate estimates, we take the average of the confidence estimates over the $K$ estimates as the final confidence estimate $C_{AVG}$, and choose the most common predicted label as the answer $\mathcal{A}_{FINAL}=\mbox{Mode}(\mathcal{A}_{1},\dots,\mathcal{A}_{K})$.
As described in Table 1, we propose to incorporate four CEMs: Base, Chain-of-Thought (CoT), Multi-Step (MS), and $K$-Averaged CoT (Self Consistency Prompting [64] or SC). The first three methods are based on deterministic sampling. SC allows us to assess attack effectiveness under more random variation. We note that MS produces far less confident responses allowing us to ascertain attack performance in lower confidence scenarios. We focus on these CEMs since they are the most straightforward to attain and directly capture the model’s perception of confidence with minimal processing. Examples of prompts can be found in Appendix T.
@@
SC Think step by step, produce letter answer and confidence at the end. Sample and average multiple answers. (1) … therefore the answer is B, 80% (2) I believe … so the answer is B, 90% (3) …, the answer must be A, 95%. Final Answer B, 88%
@@ 4 Experiment Setup
We use popular Question Answering (QA) benchmark datasets that cover the generic, medical, and legal domains: (1) MedMCQA (MQA) [47] (2) TruthfulQA (TQA) [36] (3) StrategyQA (SQA) [16] We also use the AdvGLUE [61] adversarial benchmark for confidence attack agnostic results, specifically the SST-2 [57] set. Details can be found in Appendix C.
@@
We focus on two LLMs of smaller and larger scale parameter sizes: Llama-3-8B [1] and GPT-3.5-turbo [45]. We also validate the results on GPT-4o [45] and Llama-3-70B [1]. Refer to Appendix C for details.
@@
Table 2: Effectiveness of different attack methods when targeting questions in QA prompts against a range of CEMs. Second column of Avg. shows the average distance between Cf. and Adv. Cf.
@@ 5.1 Confidence Agnostic Attack
We first determine whether CEMs are robust against generic confidence-agnostic adversarial attacks by comparing the confidences generated for the original instances in SST-2 [57] and their AdvGLUE [61] versions. The results are in Table 7 in Appendix E, where the difference in confidences are minor, with at most a 5% drop. CEMs are shown to be generally robust against confidence-agnostic perturbations, even though they are effective in reducing accuracy.
@@ 5.2 Perturbation-based Attack
To assess our proposed attack framework, we examine the four main VCAs against the CEMs across the three main datasets. For each dataset we randomly select 200 examples and test attacks against both correctly and incorrectly answered ones with the average performance across the examples. The results with user queries as the target are outlined in Table 4. We can see that the attacks are effective at reducing confidence on adversarial examples, yielding greater decreases in confidence scores compared to confidence-agnostic attacks. Nevertheless, the CEMs are largely resistant to drops in confidence of more than 50%, and the range of examples that experience a drop ranges from less than 5% up to 85%. When the attacks are effective, they lead to average confidence drops of 20%, which can have a great effect on human perception of veracity (consider a confidence of 90% versus 70%). The MS CEM, with far lower average confidence levels, is significantly more susceptible to the attacks in terms of confidence reduction, and in general the scenarios with lower average confidence are more susceptible, revealing that the greatest risk is for less confident approaches. A high percentage of the attack methods lead to answer changes, often flipping 50% up to 100% of answers, showing our framework leads to success in terms of the traditional adversarial objective.
Table 3: Averaged results for system prompt and demonstration attacks across all methods and models.
VCA-TF
Adv. Cf. % Aff. #Iters ΔAff. Cf %Flp(OC) %Flp(OW)
Sys. 5.9 24.3 56.1 22.5 21.8 30.4
Demo. 4.5 27.0 22.2 15.4 27.3 43.0
@@
Sys. 6.2 25.2 70.1 22.2 22.1 32.2
Demo. 5.0 29.1 27.1 16.6 30.3 44.3
SubSwapRemove
Sys. 11.2 43.0 58.1 24.5 46.0 68.3
Demo. 9.0 43.3 34.1 17.6 54.6 80.2
Typos
Sys. 5.1 40.6 24.5 14.2 47.0 63.2
Demo. 1.3 9.3 25.1 14.8 13.8 32.6
In Table 3 we present the averaged results across all datasets and models when using system prompts and demonstrations as attack vectors, focusing on the most robust CEMs in the Base and CoT methods (For full results, refer to Table 8 in Appendix F). System prompts and demonstrations are more vulnerable to attack, with system prompt attacks being more effective at reducing confidence of the two, while demonstration-based attacks lead to higher rates of answer changes.
Table 4: Sub-sample of attacks using Llama-3-70B and GPT-4o (averaged across all datasets and CEMs).
SubSwapRemove
Δ Cf % Aff. Δ Aff. Cf %Flip(OC) %Flips(OW)
Llama-3-70B Avg. 4.7 32.5 13.2 26.6 61.5
GPT-4o Avg. 4.8 37.9 11.1 13.3 32.5
VCA-TB
Llama-3-70B Avg. 4.2 27.3 13.7 13.0 29.5
GPT-4o Avg. 6.6 46.7 11.9 25.0 58.4
Results on GPT-4 and Llama-3-70B.
Due to the high cost of adversarial attacks, we design an experiment using the SOTA models, GPT-4 and Llama-3-70B, and run them over SSR and TB, covering the most effective of the different attack classes. The results (targeting user queries) are in Table 4 (Full Results in Table 9), drawing the same conclusions as those in Table 4.
Table 5: The effectiveness of ConfidenceTriggers, ConfidenceTriggers-AutoDAN, and a random tokens baseline are in reducing downstream confidence scores. The highest averaged drops ConfidenceTriggers across all methods of confidence for a given model are in bold. The highest individual average drop for each method is underlined. Negative Δ’s show that average confidence increased post attack.
@@
Δ  Conf % Aff. ΔAff. Cf %Flp(OC) %Flp(OW) Δ Conf % Aff. ΔAff. Cf %Flp(OC) %Flp(OW) Δ Conf % Aff. ΔAff. Cf %Flp(OC) %Flp(OW)
Llama-3-8B TruthfulQA Base 0.9 17.0 5.0 13.8 40.7 3.6 18.8 19.1 16.7 33.6 0.4 16.5 1.4 0.6 15.6
@@ 5.3 Jailbreak-based Attacks
The performance of the optimized triggers (on examples not used to optimize triggers) in both the original and AutoDAN variation are shown in Table 5. Additionally, a random tokens baseline is included as a control. For this, we generate a prompt consisting of $M=20$ randomly selected confidence related words that are appended to the system prompt. The highest of confidence drops out of all tested methods (up to 29.5%) can be achieved with ConfidenceTriggers, although with inconsistency. ConfidenceTriggers-AutoDAN is less effective but can still lead to large drops in average confidence (up to 16.7%). Both approaches outperform the random tokens baseline. ConfidenceTriggers also leads to a high percentage of answer changes, although the AutoDAN variant is less effective. Nevertheless, we show that triggers from the AutoDAN variant transfer well across datasets in Appendix S.
@@ 5.4 Stability of Confidence
To understand the stability of verbal confidence, we track confidence and answers as words are removed in succession based on their importance score from user queries. This allows us to understand how confidence and answers change as the most critical words are removed, and when the input is completely stripped of its original contents. Figure 2 shows that the average confidence scores fluctuate inconsistently as tokens are removed. Even with most of the sequence missing, the difference remains below 15% compared to the original. This also demonstrates that to most effectively reduce verbal confidence (such as above 30%) requires optimization. In contrast, answer changes follow a consistent pattern as they occur more often when the sequence is corrupted more. As a result, answer generation is more responsive to relevant factors in a sequence of text than confidence.
In Appendix G, we continue to study input corruptions by masking or randomizing confidence scores for each sub-step when using the MS method and observing how the LLM shifts its final predicted confidence when generating based on the corrupted input. The effects are found to be minor, with even the most extreme cases only leading to a drop of 30-40%. Consequently, we determine that generated confidence scores are highly stable with regards to serious text corruptions.
@@ 5.5 Confidence Behaviour
To understand the invariance of the tested LLMs, we provide the distribution of generated confidence scores in Appendix M, showing that LLMs rarely predict low confidences. We also determine that the most significant confidence changes occur on examples where the predicted answer changes (Appendix O), revealing how LLMs are most vulnerable when unsure of an answer. Nonetheless, in Appendix P we test attack variants prioritized on only manipulating confidence and show that overall confidence vulnerability is not strongly tied to whether the model’s answer is changed due to adversarial perturbations.
We further evaluate attacks effectiveness when: (1) using a lower threshold for semantic similarity (Appendix K), and find that increases in attack effectiveness are minor, (2) LLMs are prompted to provide confidence from a Likert Scale (Appendix I) noting that attacks affect these disparate manners of elicitation similarly, (3) using the Self-Probing CEM (Appendix J) finding that generating answers and confidence score in seperate steps brings few benefits. Nevertheless, in Appendix H, when different phrases that ask LLMs to change their behaviour in terms of confidence are added to prompts, there are very large shifts in confidence. On the whole, these results reveal that LLMs’ maintain a high level of certainty irrespective of input, and behaviour only significantly changes when explicitly instructed.
@@
Δconf O. Δconf A. %Aff. %Adv Aff. ΔAdv. Δconf O. Δconf A. %Aff. %Adv Aff. ΔAdv. Δconf O. Δconf A. %Aff. %Adv Aff. ΔAdv.
@@
Finally, we study alignment between numeric confidence scores and token log probability confidence in Appendix L, where we conclude that attacks create heightened misalignment in adversarial examples. This hurts model honesty by demonstrating that consistency between internal states and verbal outputs is degraded through minor perturbations. Amid the growing importance of calibration analysis, we observe how a notable degree of underconfidence is induced after attacks which aligns with our expectations (Appendix N). Despite generally high confidence levels, we highlight that LLM calibration does not entail honesty or lack thereof, since a model can become calibrated solely though imitation of a calibrated individual [73], hence we stress the importance of examining robustness and honesty regardless of inherent model calibration patterns.
5.6 Defences Against VCA
@@
We examine three typical categories of defences [68]: two input permutation defences (Paraphrase defence [26] and SmoothLLM [55]) and one self-processing perplexity filter defence [26]. Refer to Appendix R for details. We experiment with jailbreak-based VCAs due to their more generalizable nature and many recent works are devoted to mitigation strategies against jailbreaks. We center our analysis on ConfidenceTriggers with the AutoDAN variant, since the more naturalistic triggers offers greater difficulty to filter-based techniques compared to the main variant. Llama-3-8B serves as the target model. We use the same setup as for Table 5. (Results on GPT-3.5 are in Appendix Q).
@@
Table 6 presents the results for the two input permutation defences. We show the difference in average confidence provided by the LLM between the original set of data and those with the permuted prompts without any trigger phrase (Δ conf O.) and with it (Δ conf A.). Additionally, we note what percentage of permuted examples with originally benign (Aff.) and trigger-included prompts (Adv Aff.) had an altered generated confidence level compared to the original un-permuted version. Finally, we record the average decrease in confidence after input permutation on the trigger-included prompts (Δ Adv). Our primary findings are that these defences lead to significant changes in confidence both when the original prompt does and does not include a trigger phrase (differences up to  43%). The majority of examples have their confidence altered as a result of these defences and minimal reductions in attack effectiveness after permutation of an adversarial prompt signal that these defences are largely ineffective.
@@
Table 20 in Appendix Q presents the results of the perplexity filter and LLM-Guard, focusing on the percentage of user queries (that had confidence drop) that get filtered and which percentage of prompts with no trigger that would get filtered. In most cases for both methods the majority of these examples do not get filtered.

### v3

@@ 3.3.2 ConfidenceTriggers-AutoDAN
As an alternative method for more naturalistic triggers, we propose a variant of ConfidenceTriggers based on AutoDAN [liu2024autodan] (ConfidenceTriggers-AutoDAN). AutoDAN is an algorithm for automatically generating stealthy jailbreak prompts through a hierarchical genetic algorithm. Featuring the same structure as ConfidenceTriggers, hierarchical perturbations are produced each generation at the sentence-level through synonym replacement based on scores from a momentum word score dictionary, and at the paragraph-level using GPT4 [OpenAI2024] to rephrase pairs of parent prompts for Cross-Over and Mutation. Full details and the algorithm are provided in Appendix D.
@@ 3.4 Design of Confidence Elicitation
SC Think step by step, produce letter answer and confidence at the end. Sample and average multiple answers. (1) … therefore the answer is B, 80% (2) I believe … so the answer is B, 90% (3) …, the answer must be B, 95%. Final Answer B, 88.33%
Our confidence elicitation methods are based on the framework proposed in [xiong2024can], where CEMs primarily consist of a prompting strategy (i.e., task instruction) for generating verbalized numeric confidence scores. CEMs can also include sampling strategies for generating multiple LLM responses, which are then aggregated to ideally produce a more representative confidence. We use the self-random sampling strategy, which samples $K$ different output sequences $\{\bm{\mathcal{Y}}_{1},\bm{\mathcal{Y}}_{2},\dots,\bm{\mathcal{Y}}_{K}\}$ using a specific temperature parameter, from which confidence estimates $\{\mathcal{C}_{1},\dots,\mathcal{C}_{K}\}$ are derived. To aggregate these disparate estimates, we take the average of the confidence estimates over the $K$ estimates as the final confidence estimate $C_{AVG}$, and choose the most common predicted label as the answer $\mathcal{A}_{FINAL}=\mbox{Mode}(\mathcal{A}_{1},\dots,\mathcal{A}_{K})$.
As described in Table 1, we propose to incorporate four main CEMs: Base, Chain-of-Thought (CoT), Multi-Step (MS), and Self Consistency Prompting (SC) [wang2023selfconsistency]. The first three methods are based on deterministic sampling. SC allows us to assess attack effectiveness under more random variation by averaging confidence across multiple sampled responses. Note that we utilize an additional form of SC called SC-CA (Self Consistency - Complete Average) in certain experiments, most notably for developing ConfidenceTriggers, since utilizing absolute model confidence is beneficial for optimizing the algorithm. While SC averages the confidence scores of only the majority voted answer out of K predicted samples, SC-CA is concerned with absolute confidence and hence averages the confidence scores of all K predicted samples irrespective of the majority voted class. This is useful in cases where priority is on confidence irrespective of label (please refer to Appendix B for the formulation of SC-CA). We also note that MS produces far less confident responses allowing us to ascertain attack performance in lower confidence scenarios. We focus on these CEMs since they are the most straightforward to attain and directly capture the model’s perception of confidence with minimal processing. Examples of prompts can be found in Appendix W.
@@ 4 Experiment Setup
We use popular Question Answering (QA) benchmark datasets that cover the generic, medical, and legal domains: (1) MedMCQA (MQA) [pmlr-v174-pal22a] (2) TruthfulQA (TQA) [lin2022truthfulqa] (3) StrategyQA (SQA) [geva2021strategyqa] We also use the AdvGLUE [wang2021adversarial] adversarial benchmark for confidence attack agnostic results, specifically the SST-2 [socher-etal-2013-recursive] set. Details can be found in Appendix C.
@@
We focus on two LLMs of smaller and larger scale parameter sizes: Llama-3-8B [llama3modelcard] and GPT-3.5-turbo [OpenAI2024]. We also validate the results on GPT-4o [OpenAI2024] and Llama-3-70B [llama3modelcard]. Refer to Appendix C for details.
@@
Table 2: Effectiveness of different attack methods when targeting questions in QA prompts against a range of CEMs. Second column of Avg. shows the average distance between Cf. and Adv. Cf. Underlined are the most effective CEM results for each respective model/dataset/VCA method combination. In general, we can observe significant differences in Δ Aff. Cf. post attack (up to 70%) and high percentages of affected samples (up to 85%) from VCAs.
@@ 5.1 Confidence Agnostic Attack
We first determine whether CEMs are robust against generic confidence-agnostic adversarial attacks by comparing the confidences generated for the original instances in SST-2 [socher-etal-2013-recursive] and their AdvGLUE [wang2021adversarial] versions. The results are in Table 8 in Appendix E, where the difference in confidences are minor, with at most a 5% drop. CEMs are shown to be generally robust against confidence-agnostic perturbations, even though they are effective in reducing accuracy.
Figure 2: Averaged system prompt (Sys.) and demonstration attack (Demo.) performance across all models and the Base and CoT CEMs combined. Blue area over bars shows relative gains over equivalent query attacks while red show losses. We observe a high level of affected samples and significant average differences in confidence.
@@ 5.2 Perturbation-based Attack
To assess our proposed attack framework, we examine the four main VCAs against all of the CEMs across the three main datasets. For each dataset we randomly select 200 examples and test attacks against both correctly and incorrectly answered ones with the average performance across the examples. The results with user queries as the target are outlined in Table 2. We can see that the attacks are effective at reducing confidence on adversarial examples, yielding greater decreases in confidence scores compared to confidence-agnostic attacks. Nevertheless, the CEMs are largely resistant to drops in confidence of more than 50%, and the range of examples that experience a drop ranges from less than 5% up to 85%. When the attacks are effective, they lead to average confidence drops of 20%, which can have a great effect on human perception of veracity (consider a confidence of 90% versus 70%). The MS CEM, with far lower average confidence levels, is significantly more susceptible to the attacks in terms of confidence reduction, and in general the scenarios with lower average confidence are more susceptible, revealing that the greatest risk is for less confident approaches. A high percentage of the attack methods lead to answer changes, often flipping 50% up to 100% of answers, showing our framework leads to success in terms of the traditional adversarial objective.
In Figure 2 we present the averaged results across all datasets and models when using system prompts and demonstrations as attack vectors, focusing on the most robust CEMs in the Base and CoT methods (for full results, refer to Table 9 in Appendix F). System prompts and demonstrations are often more vulnerable to VCAs compared to queries, with the former being more effective at reducing confidence of the two, while demonstration-based attacks lead to higher rates of answer changes.
Table 3: Sub-sample of attacks using Llama-3-70B and GPT-4o (averaged across all datasets and CEMs). Best performance is underlined.
SubSwapRemove
Δ Cf. % Aff. Δ Aff. Cf. %Flp(OC) %Flp(OW)
Llama-3-70B Avg. 4.3 32.4 12.4 29.5 56.3
GPT-4o Avg. 5.7 48.0 11.4 27.4 53.6
@@
Llama-3-70B Avg. 3.8 26.6 12.8 14.9 28.4
GPT-4o Avg. 4.8 39.4 10.9 16.3 31.2
Results on GPT-4o and Llama-3-70B.
Due to the high cost of adversarial attacks, we design an experiment using the SOTA models, GPT-4o and Llama-3-70B, and run them over SSR and TB, covering the most effective of the different attack classes. The results (targeting user queries) are in Table 3 (full results in Table 10), confirming the findings in Table 2. In Table 11 we further test an additional three recently released models in Llama-3.1-8B [llama3modelcard], Llama-3.3-70B [llama3modelcard], and o3-mini [OpenAI2024] to show consistent VCA performance across recent architectures.
Table 4: The effectiveness of ConfidenceTriggers, ConfidenceTriggers-AutoDAN, and a random tokens baseline are in reducing downstream confidence scores. The highest averaged drops ConfidenceTriggers across all methods of confidence for a given model are in bold. The highest performances for each metric are underlined. Negative Δ’s show that average confidence increased post attack. We observe large Δ Cf. induced by optimizing a single set of tokens.
@@
Δ  Cf. % Aff. ΔAff. Cf. %Flp(OC) %Flp(OW) Δ  Cf. % Aff. ΔAff. Cf. %Flp(OC) %Flp(OW) Δ Cf. % Aff. ΔAff. Cf. %Flp(OC) %Flp(OW)
Llama-3-8B TQA Base 0.9 17.0 5.0 13.8 40.7 3.6 18.8 19.1 16.7 33.6 0.4 16.5 1.4 0.6 15.6
@@ 5.3 Jailbreak-based Attacks
The performance of the optimized triggers (on examples not used to optimize triggers) in both the original and AutoDAN variation are shown in Table 4. Additionally, a random tokens baseline is included as a control. For this, we generate a prompt consisting of $M=20$ randomly selected confidence related words that are appended to the system prompt. Very high confidence drops (up to 29.5%) can be achieved with ConfidenceTriggers, although with inconsistency. ConfidenceTriggers-AutoDAN is less effective but can still lead to large $\Delta$ confidence (up to 16.7%). Both approaches outperform the random tokens baseline. ConfidenceTriggers also leads to a high percentage of answer changes, although the AutoDAN variant is less effective. Nevertheless, we show that triggers from the AutoDAN variant transfer across datasets in Appendix U.
@@ 5.4 Stability of Confidence
To understand the stability of verbal confidence, we track confidence and answers as words are removed in succession based on their importance score from user queries. This allows us to understand how confidence and answers change as the most critical words are removed, and when the input is completely stripped of its original contents. Figure 3 shows that the average confidence scores fluctuate inconsistently as tokens are removed. Even with most of the sequence missing, the difference largely remains below 15% compared to the original. This also demonstrates that reducing verbal confidence effectively requires specific optimization rather than degrading the input as much as possible. In contrast, answer changes follow a consistent pattern as they occur more often when the sequence is corrupted more. As a result, answer generation is more responsive to relevant factors in a sequence of text than confidence.
In Appendix G, we continue to study input corruptions by masking or randomizing confidence scores for each sub-step when using the MS method and observing how the LLM shifts its final predicted confidence when generating based on the corrupted input. The effects are found to be minor, with even the most extreme cases only leading to a drop of 30-40%. Consequently, we determine that generated confidence scores are stable with regards to serious text corruptions.
@@ 5.5 Confidence Behaviour
To understand the invariance of the tested LLMs, we provide the distribution of generated confidence scores in Appendix M, showing that LLMs rarely predict low confidences. We also determine that the most significant confidence changes occur on examples where the predicted answer changes (Appendix P), revealing how LLMs’ confidence is most vulnerable when they appear to be unsure of an answer. Nonetheless, in Appendix Q we test attack variants prioritized on only manipulating confidence and show that overall confidence vulnerability is not strongly tied to whether the model’s answer is ultimately changed due to adversarial perturbations.
We further evaluate attacks effectiveness when: (1) using a lower threshold for semantic similarity (Appendix K), and find that increases in attack effectiveness are minor, (2) LLMs are prompted to provide phrasal confidence from a Likert Scale (Appendix I) noting that attacks affect these disparate manners of confidence elicitation similarly, (3) using the Self-Probing CEM (Appendix J) we find that generating answers and confidence score in separate steps brings few benefits. Moreover, in Appendix H, when different phrases that ask LLMs to change their behaviour in terms of confidence are added to prompts, there are very large shifts in confidence. On the whole, these results reveal that LLMs’ maintain a high level of certainty irrespective of input, and behaviour primarily changes when explicitly instructed, though not often in ways that are expected.
Finally, we study alignment between numeric confidence scores and token log probability confidence in Appendix L, where we conclude that attacks create heightened misalignment in adversarial examples. This hurts model honesty by demonstrating that consistency between internal states and verbal outputs is degraded through minor perturbations. Amid the growing importance of calibration analysis, we observe how a notable degree of underconfidence is induced after attacks which aligns with our expectations (Appendix N). Despite generally high confidence levels, we highlight that LLM calibration does not entail honesty or lack thereof, since a model can become calibrated solely though imitation of a calibrated individual [zhou-etal-2023-navigating], hence we stress the importance of examining robustness and honesty regardless of inherent model calibration patterns. Nonetheless, in Appendix O we validate the calibration of our CEMs by presenting Brier scores [VERIFICATIONOFFORECASTSEXPRESSEDINTERMSOFPROBABILITY] across our results. The scores produced are reasonable (below those seen in works such as [galil2021disrupting]), and in most cases increase post attack showing that miscalibration is generally induced through VCAs. Lastly, we also compare AUROC scores prior to and after VCAs in Appendix V, where we find that AUROC decreases after attacks, confirming that VCAs are effective at degrading both confidence and predictive performance.
5.6 Defences Against Jailbreak-based VCAs
@@
Δ Cf. O. Δ Cf. A. %Aff. %Adv. Aff. Δ Adv. Δ Cf. O. Δ Cf. A. %Aff. %Adv. Aff. Δ Adv. Δ Cf. O. Δ Cf. A. %Aff. %Adv. Aff. Δ Adv.
@@
Jailbreak-based VCAs present a threat as although many detection systems for detecting and filtering general unsafe behaviour exist, none exist for detecting whether jailbreaking prompts are designed to influence confidence. To highlight the limitations in existing methods, we examine three typical categories of defences [xu-etal-2024-comprehensive]: two input permutation defences (Paraphrase defence [jain2023baselinedefensesadversarialattacks] and SmoothLLM [robey2023smoothllm]) and one self-processing perplexity filter defence [jain2023baselinedefensesadversarialattacks]. Refer to Appendix S for details. We experiment with jailbreak-based VCAs due to their more generalizable nature and many recent works are devoted to mitigation strategies against jailbreaks. We center our analysis on ConfidenceTriggers- AutoDAN, since the more naturalistic triggers offers greater difficulty to filter-based techniques compared to the main variant. Llama-3-8B serves as the target model. We use the same setup as for Table 4. Results on GPT-3.5 are in Appendix R.
@@
Table 5 presents the results for the two input permutation defences. We show the difference in average confidence provided by the LLM between the original set of data and those with the permuted prompts without any trigger phrase (Δ Cf. O.) and with it (Δ Cf. A.). Additionally, we note what percentage of permuted examples with originally benign (% Aff.) and trigger-included prompts (% Adv. Aff.) had an altered generated confidence level compared to the original un-permuted version. Finally, we record the average decrease in confidence after input permutation on the trigger-included prompts (Δ Adv.). Our findings are that these defences lead to significant changes in confidence with both the inclusion and exclusion of a trigger phrase (differences up to 43%). With the majority of examples having their confidence unnecessarily altered from these defences, along with minimal decreases in attack effectiveness after adversarial prompt permutation, we deduce these defences are ineffective.
@@
Table 23 in Appendix R presents the results of the perplexity filter and LLM-Guard [goyal2024llmguardguardingunsafellm], focusing on the percentage of user queries that experienced a confidence drop post VCA that get filtered, and which percentage of prompts with no trigger that would get filtered. In most cases for both methods the majority of these examples do not get filtered.
5.7 Perturbation-based VCA Defence Challenges
We detail how utilizing common defences against perturbation-based VCAs presents a major challenge. To illustrate the difficulty in applying a perplexity filter defence, in Table 6 we compare the average and standard deviation of the perplexity across normal, adversarial (generated using VCA-TF and VCA-TB), and a typical set of human generated text (500 samples from Twitter Sentiment140 [go2009twitter]) using Llama-3-8B on TQA. Through this, we show that despite the perplexity of perturbed inputs being higher than the base tested dataset, it still has a notably lower perplexity compared to real-world text obtained from social media which represents a typical range of informal and unrefined language that models can expect as input in many domains. This illustrates the difficulty of setting an effective perplexity filter without making strict assumptions about the input data distribution, as perplexity is not elevated beyond an unreasonable level.
Table 6: Using Llama-3-8B on TQA data, the average and standard deviation of the perplexity (Perp.) across original (O.), adversarial (Adv.), and data from Sentiment140 is presented.
O. Adv. Sentiment140
Avg. Std. Avg. Std. Avg. Std.
Perp. 70.3 84.1 350.1 640.2 456.7 1043.4
Given the lack of explicit mentions of confidence in perturbed adversarial examples, we demonstrate in Table 27 that by using a simple filter from prompting GPT-4.1 (detailed in Appendix T) to flag any inputs trying to manipulate confidence scores, it is possible to detect all direct confidence statements but none without them. This shows that while directly asking an LLM for a desired confidence level is an effective attack, it is very easily detectable compared to targeting confidence using indirect means like minor word or character level modifications, hence why they are the focus of our study.